sorry if this is not the right topic; I don't know how to identify the type of problem. If 3x^2-2x+7=0, then (x- (1/3))^2= ? Answer and steps to get there would be much appreciated, thanks so much, Mike
Answer by dabanfield(803)

sorry if this is not the right topic; I don't know how to identify the type of problem. If 3x^2-2x+7=0, then (x- (1/3))^2= ? Answer and steps to get there would be much appreciated, thanks so much, Mike
3x^2 - 2x + 7 = 0
You can use the quadratic formaula to solve this quadratic equation for x. In this case a = 3, b=-2 and c=7.